OpenStudy (anonymous):

Person A speaks the truth 88% of the time. The probability of person B speaking the truth on an occasion that person A also speaks the truth is 43%. What is the probability that person A speaks the truth, but person B lies?

OpenStudy (anonymous):

my guess is \(.88\times .57\) but i could be wrong

OpenStudy (anonymous):

we can try it with numbers lets say we do it 10,000 times then person A speaks the truth 8800 times and of those 8800 43% of the time person B also speaks the truth so 3784 times the rest of them or 5016 times person A speaks the truth and person B lies

OpenStudy (anonymous):

so the probability is \(\frac{5016}{10000}=.5016\) now lets try \[.88\times .57=.5016\] yup seems to work

OpenStudy (anonymous):

yup i got that

OpenStudy (anonymous):

thank you you were a big help

OpenStudy (anonymous):

yw